OpenStudy (anonymous):

4 1/3 * 2 3/5 Multiply A. 1 2/3 B. 1 4/5 C. 2 1/5 D. 2 3/5

OpenStudy (anonymous):

Please help

OpenStudy (dan815):

are they all mixed fractions

OpenStudy (anonymous):

Multiply

OpenStudy (dan815):

11 4/5

OpenStudy (dan815):

but please learn!! dont take my word for it

OpenStudy (dan815):

make them all into common fractions

OpenStudy (anonymous):

11/4/5 reduced to lowest terms

OpenStudy (dan815):

4 1/3 = (4*3-1) /3 = 13/3 2 3/5 = (5*2+3)/5 = 13/5 13/3 * 13/5 = 159/5 =(155 + 4)/5 = 155 /5 +4/5 = 11+ 4/5

OpenStudy (dan815):

see ya drakula

OpenStudy (anonymous):

How did you know I'm drakula????

OpenStudy (dan815):

^^

OpenStudy (anonymous):

Hahahaha..

OpenStudy (dan815):

:D

OpenStudy (anonymous):

Dannnnnnnnn!